OpenStudy (anonymous):

z²-6z-27=0 Find the solutions of the equation. (I need a quick walk though)

OpenStudy (anonymous):

@Directrix do you have time?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

i didnt think that would give me the answer

OpenStudy (anonymous):

yeah i think it does factor

OpenStudy (anonymous):

how? im not that good at factoring

Directrix (directrix):

Look for two numbers that multiply to -27 and add to -6. (z - ?) * (z + ?) @JaneDoe100

OpenStudy (anonymous):

3 -9

OpenStudy (anonymous):

@Directrix is that all i need to do?

Directrix (directrix):

Those are not the final factors of z²-6z-27=0. z²-6z-27=0 (z - 9) * (z + 3) = 0 Use the Zero Product Property to write: (z - 9) = 0 or (z + 3) = 0 Solve each of those two equations for z. Those two values will be your solutions.

OpenStudy (anonymous):

9 and -3

Directrix (directrix):

That is what I got.